length of a rectangle is 8 in less than twice the width. the area is 90 in squared find the width
Answer by checkley77(12844)

L=2W-8
AREA=LW
90=(2W-8)W
90=2W^2-8W
2W^2-8W-90=0
2(W^2-4W-45)=0
2(W-9)(W+5)=0
W-9=0
W=9 ANS. FOR THE WIDTH.
L=2*9-8
L=18-8
L=10 ANS. FOR THE LENGTH.
PROOF:
90=10*9
90=90